Before planting her first tomato plants, Candy Chemist must prepare the soil for the optimum pH, which for tomatoes, is about 6.40. In her chemical cabinet, Candy finds 5.00 L bottles labeled 1.50 M club soda, 1.50 protos pion, and 1.50 M caustic soda. From her general chemistry studies, she knows club soda is simply carbonic acid (H2CO3), which is a diprotic acid with Ka1= 4.22x10-11. From her Greek, biology and chemistry, she knows that protos pion is first-fat acid, which was the unnpleasent smell of body odor, is better known as propionic acid (C2H5COOH), and is a monoprotic acid with Ka= 1.34x10-5. Candy determine that 1.00 L of the appropriate buffer solution at pH 6.40 is required to treat the soil.

Find the volume of 1.50 M caustic soda (NaOH) required to partially neutralize the appropriate acid and prepare 1.00 L of the buffer solution. You may assume all volumes are additive.
Answer #1

for a pH= 6.40

we clearly need propionic acid

pKa = -log(Ka) = -log(1.34*10^-5) = 4.87289

pH = pKa + log(A-/HA)

6.40 = 4.87289 + log(A-/HA)

A-/HA = 10^(6.40-4.87289 ) = 33.6596

assume we use 1 mmol of propionc acid

HA = 1 mmol

so we need 33.6596 mmol of NaOH; 33.6596 +1 = 34.6596 mmol of HA

V = mmol/M = 34.6596 /1.5 = 23.1064 mL

V = mmol/M = 33.6596 /1.5 = 22.439 mL

now

note that in order to maximize the ratio,

23.1064 *20 = 462 mL

22.439 *20 = 448 mL

for total V= 1000 mL, add water up to finish line V = 1000 ml